Sita expect her future earning to bw woeth es 100 if she falls ill her expect earning will be ra 25 .there ia bwlowf that aheay fall ill wit probabilty og 2/3 while the remIning in good health is 1/3 let her utility functio given by u(y) = y1/3 .supooae that insurance company offers to fully ensure sit against loss of earning cauaed by illness aganist actuarally fair premium (a) will sita L accept insurance explain (b) what ia max amt .that sita would pay fr insurance
1
Expert's answer
2017-06-09T10:58:12-0400
Future earnings - Rs. 100, p = 1/3
If she falls ill - Rs. 25, p = 2/3
U(y) = y1/2
(a) Sita may accept the insurance, if the probability of illness is too high and the insurance payment is affordable. As p = 2/3, Sita will accept the insurance.
(b) The maximum amount that Sita would pay for the insurance can be calculated according to probabilities:
i(max) = 100*1/3 + 25*2/3 = 33.33 + 16.67 = 50.